PLEASE HELP ME SOLVE:
BOX HAS TOTAL OF 44 COINS, CONSISTING OF PENNIES, DIMES AND QUARTERS. THE NUMBER OF PENNIES IS EQUAL TO THE NUMBER OF DIMES, AND THE TOTAL VALUE IS $4.37 HOW MANY OF EACH DENOMINATION OF COIN DOES HE HAVE.

Answer by checkley71(8403)

44=2x+q & 4.37=.01x+.10x+.25q or q=44-2x or 4.37=.01x+.10x+.25(44-2x) or
4.37=.0x+.10x+11.00-.50x or 4.37-11.00=.01x+.10x-.50x or -6.63=-.39x or
6.63/.39=x or x=17 thus there are 17 pennies, 17 dimes & 10 quarters
proof 17+17=10=44 & .01*17+.10*17+.25*10=4.37 .17+1.70+2.50=4.37 or 4.37=4.37